OpenStudy (anonymous):

What is the equation of the line that passes through the point (2, 5) and has a slope of 2?

OpenStudy (the_fizicx99):

Plug in those given values in: \[y-y1=M(x-x1)\]

OpenStudy (the_fizicx99):

I'll start it for you so you have an idea of what it looks like, \[y-5=2(x-2)\]

OpenStudy (anonymous):

y = 2x + 1?

OpenStudy (the_fizicx99):

y - 5 = 2x - 4 +5 +5 y = 2x + 1 Yes you're right! good job :3

OpenStudy (anonymous):

Thanks

OpenStudy (the_fizicx99):

yw